OpenStudy (anonymous):

inequalitys ... how do i get rid of 29? 21x=3x-29

OpenStudy (the_fizicx99):

Combine the like terms, subtract 3x from both sides and divide to isolate x. 21x = 3x - 29 -3x -3x 18x = -29 /18 /18 x = ?

OpenStudy (anonymous):

suntract the 3X then you'll get 18X= -29 then divide by 18 and thats your answer

OpenStudy (anonymous):

1.61

OpenStudy (anonymous):

-1.61

OpenStudy (the_fizicx99):

You're dividing by a negative. One of them is positive, a negative and positive results in negative. -1.61